Research examining the effects of preschool child care has found that children who spent time in...

  1. Research examining the effects of preschool child care has found that children who spent time in day care, especially high-quality day care, perform better on math and language tests than children who say home with their mothers (Broberg, Wessels, Lamb, & Hwang, 1997). In a typical study, a researcher obtains a sample of n = 10 children who attended day care before starting school. Then children are given a standardized math test for which the population mean is μ = 52.

The scores for the sample are as follows:

               53,   57,   61,   49,   52,   56,   58,   62,   51,   56

On the basis of this sample, children with a history of preschool day care significantly different from the general population? Use a two-tailed test with α = .05.

  1. State the null hypothesis in words and in a statistical form
  1. State the alternative hypothesis in words and a statistical form

  1. Compute the appropriate statistic to test the hypotheses. Sketch the distribution with the estimated standard error and locate the critical region(s) with the critical value(s) (6). [When you compute SS, you might want to use the definitional formula as it might be easier to compute SS this way].
  1. State your statistical decision
  1. Compute Cohen’s d. Interpret what this d really means in this context

f.    Compute 95% CI

                       

g. What is your conclusion? Interpret the result. Don’t forget to include statistical information as well (e.g., t-score, df, α, Cohen’s d)
Answer #1

Σx = 555 Σx2 = 30965 ΣΧ Mean i n 53 + 57 + 61 + 49 + 52 + 56 + 58 + 62 + 51 + 56 Π 10 555 10 = 55.5 (Σx) 2 Σχ2 n Sample StandThe provided sample mean is I = 55.5 and the sample standard deviation is s = 4.2492, and the sample size is n = 10. (1) Null(4) Decision about the null hypothesis Since it is observed that t| = 2.605 > te= 2.262, it is then concluded that the null hThus there is sufficient evidence to conclude that children with history of preschool day care significantly different than true population.

The provided sample mean is X = 55.5 and the sample standard deviation is s = 4.2492. The size of the sample is n = 10 and ththus we are 95% confident that the mean maths score of children who spend time in day care lies between 52.46 to 58.54

I have given my best to solve your problem. Please like the answer if you are satisfied with it. ?